A living 120-pound woman contains about 6 x 10 grams of carbon-14. How many grams of carbon-14 would remain after 17100 years? What has happened to the rest of the carbon-14? (Assume carbon-14 is a beta emitter with a half-life of 5700 years.)
1
Expert's answer
2020-09-16T07:52:13-0400

Half-life (t1/2)is the time it can take for half the mass of a radio-active isotope to undergo decay.

Since for carbon-14 this time is given as 5700 years; then it follows that after this period, any mass of carbon-14 would have decayed, and only half of it will remain.

In simple terms:

After 17100 years, the number of half-lifes for carbon-14 are:

"=\dfrac{17100}{5700} = 3 half-lifes"


This then means;

For the first half-life; "\dfrac{1}{2}x60g =30g" (30g will decay and 30g remain)


For the second, "\dfrac{1}{2}x30 = 15g" (15g decay and 15 remain)


For the third; "\dfrac{1}{2}x15 = 7.5g" (7.5g decay and 7.5 remain)

Thus after 17100 years, 7.5 g of the initial 60g will remain.


When carbon-14 undergoes decay, it is called a beta decay because it emits an electron and an electron antineutrino.


14C "\to" 14N +e- + ve-

One of the neutrons become a proton and the carbon-14 decays into a stable(non-radio-active) isotope nitrogen-14